# Assessing Web Search Credibility and Response Groundedness in Chat Assistants

### v1

@@ Abstract
Chat assistants increasingly integrate web search functionality, enabling them to retrieve and cite external sources. While this promises more reliable answers, it also raises the risk of amplifying misinformation from low-credibility sources. In this paper, we introduce a novel methodology for evaluating assistants’ web search behavior, focusing on source credibility and the groundedness of responses with respect to cited sources. Using 100 claims across five misinformation-prone topics, we assess GPT-4o, GPT-5, Perplexity, and Qwen Chat. Our findings reveal differences between the assistants, with Perplexity achieving the highest source credibility, whereas GPT-4o exhibits elevated citation of non-credibility sources on sensitive topics. This work provides the first systematic comparison of commonly used chat assistants for fact-checking behavior, offering a foundation for evaluating AI systems in high-stakes information environments.
@@ 1 Introduction
Chat assistants powered by large language models (LLMs) are increasingly used for information seeking Chatterji et al. (2025). With integrated web search, they can retrieve and cite relevant sources instead of relying only on internal knowledge. This offers opportunities for assistants to ground their answers in up-to-date evidence. However, this functionality also raises the critical challenge that retrieved evidence may come from disinformation sources American Sunlight Project (2025). When chat assistants cite unreliable sources or present fabricated information with high confidence, they risk amplifying misinformation instead of mitigating it Zellers et al. (2019); Vykopal et al. (2024).
These risks are pressing in domains, such as health Waszak et al. (2018), climate change van der Linden et al. (2017), or political discourse Kansaon et al. (2025), where misinformation spreads quickly and has a serious societal impact. Chat assistants retrieve dynamically from the open web, making their responses sensitive to content quality, query framing, and user beliefs. Moreover, there is growing concern that disinformation actors deliberately flood the web with propaganda to influence what AI systems retrieve: It has been shown that Russian propaganda mechanisms have successfully seeded Russian disinformation into Western AI systems Sadeghi and Blachez (2025). Yet, no prior work has explored how assistants retrieve, cite, and ground responses with web search. Understanding how web-search-enabled assistants behave when tasked with fact verification is, therefore, critical for assessing their reliability.
@@
Our methodology, shown in Figure 1, consists of three steps. The first step includes data collection (Section 3). To capture how assistants handle misinformation-prone content, we curated a list of claims spanning five topics. Each claim is tested against the assistants by modeling two distinct user roles: a fact-checker, seeking to verify the claim, and a claim believer, seeking confirmation of false information. This reflects real-world usage, where assistants face both skeptical and credulous users, and it allows us to observe how framing shapes web search and the response. This distinction is important since prior work shows that queries with false presuppositions make AI systems more likely to accept misinformation Sieker et al. (2025).
@@
Our contributions are twofold: (1) we introduce a novel methodology for evaluating web-search-enabled chat assistants that integrates credibility assessment and groundedness analysis; [^1] Code and data are available here: https://github.com/kinit-sk/web-search-analysis (2) we present the first systematic study of chat assistants’ web search functionality, showing what kinds of sources they cite, how credible these sources are, and to what extent their responses are grounded in reliable evidence.
@@ 2.2 Web Search Analysis
As chat assistants increasingly rely on external web sources, it is critical to evaluate not just what content they generate, but which sources they use, how they cite them, and how credible these sources are. Recent studies reveal large gaps in how search-enabled LLMs attribute the web content they consume. Yang (2025) examines over 65K responses across several providers, finding that news citations heavily concentrate among a few outlets, and that there is a systematic political leaning (liberal bias) in cited sources. Meanwhile, Strauss et al. (2025) reports "attribution gaps", e.g., many responses are generated without fetching any external pages, or fetch many but cite few, showing that retrieval behavior does not translate into citation or grounding.
@@ 2.3 Factuality Analysis
Early efforts such as FActScore Min et al. (2023) introduced the idea of decomposing LLMs’ outputs into atomic facts and verifying them against Wikipedia. Building on this, Factcheck-Bench Wang et al. (2024a) proposed a multi-stage pipeline for decomposition, decontextualization, retrieval, and stance detection, enabling fine-grained evaluation across claims and sentences. On the other hand, Wei et al. (2024) benchmarked long-form factuality of LLMs using a multi-step reasoning process with Google Search, while determining whether a fact is supported by the evidence. More recently, VERIFY Fatahi Bayat et al. (2025) unified these approaches into an evaluation framework, retrieving web evidence and classifying claims as supported, unsupported or undecidable, with strong correlation with human judgments.
@@ 3.3 Chat Assistant Providers
To study how current chat assistants leverage web search and employ the found sources to support their responses, we evaluated three chat providers: ChatGPT, Perplexity, and Qwen Chat. We selected available chat assistants with search functionality enabled at the time of our research. We chose to interact with the assistants through their web interfaces rather than APIs. This design choice reflects the user experience and avoids differences that may occur without using the API, since web interfaces can implement additional safety mechanisms, such as more strict filtering of disinformation. To simulate a real user, we automated interactions using Selenium [^3] https://www.selenium.dev/, mimicking real user clicks and typing.
@@ 4.2 Results
The GPT-5 model can use a thinking mode, which is automatically activated at the assistant’s discretion. In our experiments, this occurred in 10% of conversations (58 out of 600). To assess its effect, we compared conversations with and without the thinking mode.
@@ 5.2 Results
Table 4 indicates that all assistants rarely cite non-credible sources; however, the consequences of such instances differ. GPT-4o and GPT-5 occasionally cite weaker sources, especially for the Russia-Ukraine war and Local topics. Perplexity shows particularly cautious grounding behavior, maintaining high grounding while minimizing grounding on unreliable sources, similarly to the results in Section 4.2. In contrast, Qwen Chat shows a low overall reliance on non-credible material, but when such sources are used, they tend to dominate the response, with around 60% of units supported by unreliable evidence in responses that include at least one unreliable source. This behavior indicates a fragile retrieval strategy, while Qwen Chat usually avoids disinformation, but when it fails, failures result in substantial inaccuracies.
@@ 6 Discussion
Contested Topics and Claim-Believer Queries Amplify Risks.
@@ 7 Conclusion
Our study captures the state of chat assistants in the second half of 2025. Given the rapid model development and the continuous expansion of fact-checking resources, the performance we report may shift in the near future. We partially mitigated this limitation by selecting stable misinformation topics and claims. However, future replications will be needed to assess whether our findings generalize to newer systems and newer claims.
@@
We evaluated assistants without enabling additional reasoning or "thinking modes", since our goal was to assess web-search functionality as it appears to typical users. This allowed us to capture how assistants actually retrieve and ground information in practice, without relying on enhanced reasoning features that are not always visible or accessible to users. Nevertheless, automatically enabled thinking in GPT-5 demonstrated that enabling such a mode can improve source credibility and reduce exposure to disinformation. A systematic investigation of how reasoning capabilities interact with web search remains for future research.
@@
This study evaluated chat assistants using paraphrased claims derived from fact-checking organizations. While these claims represent false information, we believe that this work does not amplify the risk of spreading harmful narratives, since those claims can be found freely on the Internet. The aim of this study is to evaluate on what sources current chat assistants rely on and whether those chat assistants amplify the spread of false information in their responses. All used data were sourced from publicly available fact-checking repositories and online sources, and no private or personal information was collected.
@@ Appendix B Data & Response Collection
H17
Robert F Kennedy Jr. had won a lawsuit against big pharmaceutical companies, and the US Supreme Court had ruled that coronavirus vaccines were not vaccines
@@
C17
The aphelion phenomenon, which occurs when the Earth reaches its farthest point from the Sun, is expected to cause a drop in temperatures from May to August 2025,
with potential negative impacts on human health
@@
L17
It was found that there were attempts to intimidate KPPS carried out by certain individuals and prevented prospective voters from voting for
the Koster-Giri candidate pair in Jakarta in November 2024
@@ B.1 Chat Providers
Figure 8: Qwen Chat Interface for collecting responses.
@@ Appendix C Source Credibility Analysis
The Credibility Rate represents the proportion of the sources that are considered credible. Specifically, it defines the ratio of domains classified as very high, high, or mostly factual to the total number of classified sources. Formally:
@@
The Non-Credibility Rate captures the proportion of the cited sources that fall into the low factuality categories. Specifically, it defines the ratio of domains rated as satire, very low, or low to the total number of classified sources. Formally:
@@ D.1 Evaluation
D.2 FActScore, VERIFY and Our Method
Tables 14 and 15 compare how the three approaches characterize factual precision and hallucination across topics and user types for all chat assistants. These results do not present the performance scores, but instead, they reflect the levels of factual precision or hallucination as identified by each evaluated method.

### v2

@@ Abstract
Chat assistants increasingly integrate web search functionality, enabling them to retrieve and cite external sources. While this promises more reliable answers, it also raises the risk of amplifying misinformation from low-credibility sources. In this paper, we introduce a novel methodology for evaluating assistants’ web search behavior, focusing on source credibility and the groundedness of responses with respect to cited sources. Using 100 claims across five misinformation-prone topics, we assess GPT-4o, GPT-5, Perplexity, and Qwen Chat. Our findings reveal differences between the assistants, with Perplexity achieving the highest source credibility, whereas GPT-4o exhibits elevated citation of non-credible sources on sensitive topics. This work provides the first systematic comparison of commonly used chat assistants for fact-checking behavior, offering a foundation for evaluating AI systems in high-stakes information environments.
@@ 1 Introduction
Chat assistants powered by large language models (LLMs) are increasingly used for information seeking Chatterji et al. (2025). With integrated web search, they can retrieve and cite relevant sources instead of relying only on internal knowledge. This offers opportunities for AI assistants to ground their answers in up-to-date evidence. However, this functionality also raises the critical challenge that retrieved evidence may come from disinformation sources American Sunlight Project (2025). When chat assistants cite unreliable sources or present fabricated information with high confidence, they risk amplifying misinformation instead of mitigating it Zellers et al. (2019); Vykopal et al. (2024).
These risks are pressing in domains, such as health Waszak et al. (2018), climate change van der Linden et al. (2017), or political discourse Kansaon et al. (2025), where misinformation spreads quickly and has a serious societal impact. Chat assistants retrieve dynamically from the open web, making their responses sensitive to content quality, query framing, and user beliefs. Moreover, there is growing concern that disinformation actors deliberately flood the web with propaganda to influence what AI systems retrieve. It has been shown that Russian propaganda mechanisms have successfully seeded Russian disinformation into Western AI systems Sadeghi and Blachez (2025). Yet, no prior work has explored how assistants retrieve, cite, and ground responses with web search. Understanding how web-search-enabled assistants behave when tasked with fact verification is, therefore, critical for assessing their reliability.
@@
Our methodology, as illustrated in Figure 1, comprises three steps. The first step includes data collection (Section 3). To capture how assistants handle misinformation-prone content, we curated a list of claims spanning five topics. Each claim is tested against the assistants by modeling two distinct user roles: a fact-checker, who seeks to verify the claim, and a claim believer, who seeks confirmation of false information. This reflects real-world usage, where assistants encounter both skeptical and credulous users, and it enables us to observe how framing influences web search and the response. This distinction is important since prior work shows that queries with false presuppositions make AI systems more likely to accept misinformation Sieker et al. (2025).
@@
Our contributions are twofold. First, we propose a novel evaluation methodology for web-search-enabled chat assistants that jointly analyzes (i) the credibility of retrieved and cited sources and (ii) the groundedness of generated responses with respect to those sources. By linking groundedness to source quality, our approach reveals failure modes that factuality-only evaluation cannot detect. [^1] Code and data are available here: https://github.com/kinit-sk/web-search-analysis Second, we present the first systematic analysis of how chat assistants retrieve, cite, and ground information when using web search in fact-checking scenarios. Unlike prior studies that examine retrieval behavior or grounding in isolation, we analyze how user framing and web search jointly influence both the credibility of retrieved evidence and the reliability of grounded responses.
@@ 2.2 Web Search Analysis
As chat assistants increasingly rely on external web sources, it is crucial to evaluate not only the content they generate but also the sources they use, how they cite them, and the credibility of these sources. Recent studies reveal large gaps in how search-enabled LLMs attribute the web content they consume. Yang (2025) examines over 65K responses across several providers, finding that news citations heavily concentrate among a few outlets, and that source selection correlates with reliability and political leaning, as sources categorized as left- or center-leaning were often rated more reliable. Meanwhile, Strauss et al. (2025) reports "attribution gaps", e.g., many responses are generated without fetching any external pages, or fetch many but cite few, showing that retrieval behavior does not translate into citation or grounding.
@@ 2.3 Factuality Analysis
Early efforts such as FActScore Min et al. (2023) introduced the idea of decomposing LLMs’ outputs into atomic facts and verifying them against Wikipedia. Building on this, Factcheck-Bench Wang et al. (2024a) proposed a multi-stage pipeline for decomposition, decontextualization, retrieval, and stance detection, enabling fine-grained evaluation across claims and sentences. On the other hand, Wei et al. (2024) benchmarked long-form factuality of LLMs using a multi-step reasoning process with Google Search, while determining whether a fact is supported by the evidence. More recently, VERIFY Fatahi Bayat et al. (2025) unified these approaches into an evaluation framework, retrieving web evidence and classifying claims as supported, unsupported, or undecidable, with strong correlation with human judgments.
While these approaches advance our ability to evaluate factuality, they typically treat retrieved sources as uniformly reliable or analyze grounding independently of source quality. Similarly, prior web-search analyses such as Yang (2025) focus on retrieval patterns and citation behaviors but do not evaluate whether generated responses are grounded in credible versus non-credible evidence. In contrast, our work explicitly integrates source credibility into groundedness evaluation, enabling the identification of a critical failure mode: responses that are internally consistent with cited evidence yet rely on unreliable or disinformation sources. This allows us to examine how assistants’ retrieval choices impact the factual reliability of their outputs in ways that prior frameworks cannot capture.
@@ 3.3 Chat Assistant Providers
To study how current chat assistants leverage web search and employ the found sources to support their responses, we evaluated three chat providers: ChatGPT, Perplexity, and Qwen Chat. We selected available chat assistants with search functionality enabled at the time of our research. We chose to interact with the assistants through their web interfaces rather than APIs. This design choice reflects the user experience and avoids differences that may occur when using the API, as web interfaces can implement additional safety mechanisms, such as stricter filtering of disinformation. To simulate a real user, we automated interactions using Selenium [^3] https://www.selenium.dev/, mimicking real user clicks and typing.
@@ 4.2 Results
The GPT-5 model includes a system-controlled thinking mode, which is automatically triggered at the assistant’s discretion and cannot be explicitly disabled by the user. In our setup, the model opted to invoke this mode in 10% of the conversations (58 out of 600), typically in cases that appeared to require more complex reasoning or the synthesis of retrieved web content.
@@ 5.2 Results
Table 4 indicates that all assistants rarely cite non-credible sources; however, the impact of such citations varies. GPT-4o and GPT-5 occasionally cite weaker sources, especially for the Russia-Ukraine war and Local topics. Perplexity shows particularly cautious grounding behavior, maintaining high grounding while minimizing grounding on unreliable sources, similarly to the results in Section 4.2. In contrast, Qwen Chat generally avoids non-credible sources, but when it does cite one, a large portion of the response is affected: among responses that include at least one non-credible source, around 60% of the factual units are grounded in such sources. This behavior indicates that, while Qwen Chat usually avoids disinformation, failures can lead to substantial inaccuracies due to a fragile retrieval strategy.
@@ 6 Discussion
Contested Topics and Claim Believer Queries Amplify Risks.
@@ 7 Conclusion
Our study reflects the state of chat assistants as of the second half of 2025, and system behavior may evolve as models and fact-checking resources continue to evolve. To ensure reproducibility and comparability across assistants and topics, we primarily focused on stable misinformation claims that have persisted or repeatedly reappeared, such as narratives related to climate change or war conflicts. This choice limits our ability to fully assess assistant behavior on newly emerging claims, which is often addressed unevenly. We partially mitigate this by including some recent claims and those circulating outside English-speaking regions. We note that emerging misinformation frequently appears as variations of long-standing narratives, suggesting that stable claims still capture core patterns relevant to misinformation handling.
@@
We evaluated assistants without explicitly enabling additional reasoning or "thinking modes", as our primary objective was not to study internal reasoning strategies but to assess web-search behavior and the resulting responses as they appear to typical users. This allowed us to capture how assistants actually retrieve and ground information in practice, without relying on enhanced reasoning features that are not always visible or accessible to users. Nevertheless, in the case of GPT-5, thinking mode was automatically activated by the system in a subset of interactions and could not be disabled. Since this mode was part of the generated outputs, it could influence response quality and retrieval. Therefore, we report its effect by comparing conversations with and without the thinking mode. While automatically triggered thinking in GPT-5 was associated with improved source credibility and reduced exposure to disinformation, a systematic investigation of how reasoning capabilities interact with web search remains for future research.
@@
Groundedness Evaluation.
Our assessment of whether assistant responses were grounded in retrieved sources relied on a single LLM, Llama 3.3 70B, following the VERIFY framework Fatahi Bayat et al. (2025), which uses Llama 3 70B. This model was chosen because it correlated well with human judgments and provided consistent automated entailment scoring across atomic units. Using a single standardized model ensures comparability across assistants and avoids inter-model variability that could bias results. We acknowledge that relying on a single model may introduce biases in entailment judgments and does not capture potential differences that might arise if other models were used. However, given the scale of our evaluation, human annotation was infeasible, and our focus is on relative differences between assistants, for which a consistent automated judge is appropriate.
@@
This study evaluated chat assistants using paraphrased claims derived from fact-checking organizations. While these claims represent false information, we believe that this work does not amplify the risk of spreading harmful narratives, since those claims can be found freely on the Internet. The aim of this study is to evaluate the sources on which current chat assistants rely and whether those chat assistants amplify the spread of false information in their responses. All used data were sourced from publicly available fact-checking repositories and online sources, and no private or personal information was collected.
@@ Appendix B Data & Response Collection
H17 Robert F Kennedy Jr. had won a lawsuit against big pharmaceutical companies, and the US Supreme Court had ruled that coronavirus vaccines were not vaccines
@@
C17 The aphelion phenomenon, which occurs when the Earth reaches its farthest point from the Sun, is expected to cause a drop in temperatures from May to August 2025, with potential negative impacts on human health
@@
L17 It was found that there were attempts to intimidate KPPS carried out by certain individuals and prevented prospective voters from voting for the Koster-Giri candidate pair in Jakarta in November 2024
@@ Appendix C Source Credibility Analysis
The Credibility Rate represents the proportion of the sources that are considered credible. It defines the ratio of domains classified as very high, high, or mostly factual to the total number of classified sources. Formally:
@@
The Non-Credibility Rate captures the proportion of the cited sources that fall into the low factuality categories. It defines the ratio of domains rated as satire, very low, or low to the total number of classified sources. Formally:
@@ C.2 Additional Results
To further clarify differences observed in our experiments, we conducted statistical tests on source credibility shown in Table 11. We examined the impact of different AI assistants and user framings on source credibility. Differences across models are mostly statistically significant, indicating that the choice of assistant substantially affects the credibility of the information provided. In contrast, user framing (Fact-Checker vs Claim Believer) has a smaller but still significant effect on credibility.
CR Diff NCR Diff
Models
GPT-4o vs GPT-5 3.787 (**) 0.235 (ns)
GPT-4o vs Perplexity -11.145 (***) 1.576 (***)
GPT-4o vs Qwen Chat -4.852 (***) 1.205 (**)
GPT-5 vs Perplexity -14.932 (***) 1.340 (***)
GPT-5 vs Qwen Chat -8.639 (***) 0.969 (**)
Perplexity vs Qwen Chat 6.293 (***) -0.371 (ns)
User Type
Fact-checker vs Claim Believer 1.576 (***) 0.235 (ns)
Table 11: Pairwise comparison of source credibility across models and user framings. CR: Credibility Rate, NCR: Non-Credibility Rate. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, ns = not significant.
@@ D.1 Evaluation
D.2 Additional Results
Groundedness was evaluated using Llama 3.3 70B as an automated entailment model. Differences across AI assistants, as shown in Table 14, are mostly significant, indicating that some models produce responses that are better grounded in retrieved sources. User framing has a minor impact on groundedness, similar to source credibility.
GS Diff CG Diff NCG Diff
Models
GPT-4o vs GPT-5 -1.784 (*) 1.780 (ns) -0.270 (ns)
GPT-4o vs Perplexity -3.265 (***) -3.795 (**) 1.030 (***)
GPT-4o vs Qwen Chat 8.696 (***) 5.631 (***) 0.145 (ns)
GPT-5 vs Perplexity -1.481 (*) -5.574 (***) 1.300 (***)
GPT-5 vs Qwen Chat 10.480 (***) 3.852 (*) 0.415 (ns)
Perplexity vs Qwen Chat 11.961 (***) 9.426 (***) -0.884 (*)
User Type
Fact-Checker vs Claim Believer -1.784 (*) 8.696 (***) -3.265 (***)
Table 14: Pairwise comparison of response groundedness across models and user framings. GS: Groundedness Score, CG: Credible Groundedness, NCG: Non-Credible Groundedness. Significance levels: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, ns = not significant.
D.3 FActScore, VERIFY and Our Method
Tables 16 and 17 compare how the three approaches characterize factual precision and hallucination across topics and user types for all chat assistants. These results do not present the performance scores, but instead, they reflect the levels of factual precision or hallucination as identified by each evaluated method.
@@ D.3 FActScore, VERIFY and Our Method
Figure 4: Highlighting functionality in the GPT-4o interface.
Figure 8: Qwen Chat Interface for collecting responses.